Assume that you own a dividend-paying stock currently worth $150. You plan to sell the stock in 250 days. In order to hedge against a possible price decline, you wish to take a short position in a forward contract that expires in 250 days. The risk-free rate is 5.25% per annum (discretely compounding). Over the next 250 days, the stock will pay dividends according to the following schedule:

Days to Next Dividend

Dividends per Share ($)

30

1.25

120

1.25

210

1.25

(a) Calculate the forward price of a contract established today and expiring in 250 days.

(b) It is now 100 days since you entered the forward contract. The stock price is $115. Calculate the value of the forward contract at this point.

c) At expiration, the price of the stock is $130. Calculate the value ot the forward contract at expiration

About this question, if we have a continuos interest what is the solution ?
Answer #1
Future Value (PV) of Cash Flow:
(Cash Flow)*((1+i)^N)
i=Interest Rate=5.25%=0.0525
N=Number of years in the future
Since forward contracts needs to be settled in future date,
Forward Price will be equal to future value of Cash flows
A B=250-A C=B/360 D E=D*(1.0525^C)
Number of days Number of days to 250 Years to250days Cash flow Future value
0 250 0.694444 $150 155.4258589
30 220 0.611111 ($1.25) -1.289704419
120 130 0.361111 ($1.25) -1.273311501
210 40 0.111111 ($1.25) -1.257126947
SUM 151.6057161
(a) Forward Price of Contract $151.61
(b) Forward Price after 100 days
A B=150-A C=B/360 D E=D*(1.0525^C)
Number of days Number of days to 250 Years to250days Cash flow Future value
0 150 0.416667 $115 117.478137
20 130 0.361111 ($1.25) -1.273311501
110 40 0.111111 ($1.25) -1.257126947
SUM 114.9476985
Forward Price of Contract $114.95
.(c) Forward Price $151.61
Price at expiration (Settlement) $130
Value of Forward Contract at expiration $21.61 (151.61-130)
If we have a continuous interest :
Future value =Present value*(e^(0.0525*N)
N=number of Years in to the Future
There willnot be much difference as calculation below shows
A B=250-A C=B/360 D E=D*(e^(0.0525*C))
Number of days Number of days to 250 Years to250days Cash flow Future value
0 250 0.694444 $150              155.56966
30 220 0.611111 ($1.25) ($1.290754)
120 130 0.361111 ($1.25) ($1.273924)
210 40 0.111111 ($1.25) ($1.257313)
SUM 151.747672
Forward Price of Contract $151.75